# FactCHD: Benchmarking Fact-Conflicting Hallucination Detection

### v2

@@ Abstract
Despite their impressive generative capabilities, LLMs are hindered by fact-conflicting hallucinations in real-world applications. The accurate identification of hallucinations in texts generated by LLMs, especially in complex inferential scenarios, is a relatively unexplored area. To address this gap, we present FactCHD, a dedicated benchmark designed for the detection of fact-conflicting hallucinations from LLMs. FactCHD features a diverse dataset that spans various factuality patterns, including vanilla, multi-hop, comparison, and set operation. A distinctive element of FactCHD is its integration of fact-based evidence chains, significantly enhancing the depth of evaluating the detectors’ explanations. Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately. Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark is available at https://github.com/zjunlp/FactCHD..
@@ 1 Introduction
Large Language Models (LLMs) Zhao et al. (2023) are susceptible to generating text that, while seemingly credible, can be factually inaccurate or vague, leading to the spread of misinformation online Yin et al. (2023); Ji et al. (2023); Huang et al. (2023). This issue, referred to as fact-conflicting hallucination Zhang et al. (2023), arises from the incorporation of incorrect or obsolete knowledge into the models’ parameters and from the models’ inherent limitations in complex cognitive ability. These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information.
@@
Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
However, traditional fact verification tasks Wadden et al. (2020a, b) are not suitable for LLM-based “Query-Response” data. Moreover, existing hallucination evaluation benchmarks Li et al. (2023); Muhlgay et al. (2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. To bridge this gap, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
Datasets # Source Domains Construction Factuality Pattern Evaluation Expandability
Fever Thorne et al. (2018) Text General HG Van. Consist.
Climate-Fever Diggelmann et al. (2020) Text Climate Change HG Van. Consist.
Health-Fever Sarrouti et al. (2021) Text Health HG Van. Consist.
Sci-Fact Wadden et al. (2020a) Text Scientific HG Van. Consist.
CoVERT Mohr et al. (2022) Text COVID-19 HG Van. Consist.
TabFact Chen et al. (2020) Table General HG Van. Consist.
HoVer Jiang et al. (2020) Text General HG Mul. Consist.
FEVEROUS Aly and et al (2021) Text+Table General HG Van. Consist.
HaluEval Li et al. (2023) Text General MIX Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Van.&Mul.&Com.&Set. Fact.+Chain.
Table 1: Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
@@
  • We present FactCHD, a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
@@
  • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses. Empirical results demonstrate the effectiveness of our approach in hallucination detection.
@@ 2 Related Work
Figure 3: Domain distribution of FactCHD, where “General-p” and “General-d” denote common domains derived from Wikipedia and Wikidata, respectively.
Despite LLMs Zhao et al. (2023); Chen et al. (2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al. (2023); Ye et al. (2023); et al (2023a); Hu et al. (2023). These hallucinations, as categorized by Zhang et al. (2023), can be input-conflicting, context-conflicting, or fact-conflicting—with the latter being especially problematic due to the propagation of inaccurate factual information online. While previous studies have extensively examined hallucinations within natural language generation (NLG) for various NLP tasks Shuster et al. (2021); Creswell and Shanahan (2022); Das et al. (2023); Mallen et al. (2022), HaluEval Li et al. (2023) has emerged as a recent benchmark for assessing LLMs’ recognition of such errors. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate whether hallucinatory, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation.
@@
Our research is also related to prior works on fact verification in NLP tasks Thorne et al. (2018); Wadden et al. (2020b); Gupta et al. (2022); Dziri et al. (2022); Liang et al. (2023); Rashkin et al. (2021); Kryściński et al. (2020), expanding the understanding of factuality beyond binary judgments. The FRANK framework Pagnoni et al. (2021) offers a detailed typology of factual errors, while  Dhingra et al. (2019) explores lexical entailment in text generation. The FACTOR score by Muhlgay et al. (2023) evaluates LMs based on the likelihood of factual content. Existing benchmarks often miss the “Query-Response” context of LLMs, focusing solely on accuracy without providing explanatory rationales. Our contribution lies in: (1) presenting “Query-Response” formatted data for LLMs; (2) introducing metrics for interpretability in detecting fact-conflicting hallucinations; and (3) validating the effectiveness of Truth-Triangulator through cross-referential verification from multiple sources. The comparison with other datasets is detailed in Table 1.
@@ 4.1 Collect Realistic Data as Demonstration
Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM THUDM (2023), Alpaca, and Vicuna) to generate authentic responses. By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence. Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
@@ 4.2 Generate “Query-Response” Contexts
Figure 4: Overview of the construction process of FactCHD.
I want you to act as a data generator for a fallacy-finding task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: NON-FACTUAL
#Query#: How many children did the leader of the country, where the distributor of "The Killer Inside Me" operates, have?
#Response#: The head of government of the United States of America, John F. Kennedy, had 2 children.
#Evidence#: ["The Killer Inside Me", "distributor", "Warner Bros."], ["Warner Bros.", "owner of", "Warner Bros. Animation"], ["Warner Bros. Animation", "country", "United States of America"], ["United States of America", "head of government", "John F. Kennedy"], ["John F. Kennedy", "number of children", "4"]
#Output#: NON-FACTUAL. The answer stating that the head of government of the United States of America, John F. Kennedy, had 2 children is incorrect. "The Killer Inside Me" is distributed by Warner Bros., which is the owner of Warner Bros. Animation. Warner Bros. Animation is located in the United States of America, where the head of government was John F. Kennedy. However, John F. Kennedy had 4 children, not 2. Therefore, there is an incorrect conclusion in this query and response.
<Demonstrations>
#Golden Lable#: <Here is the right label for the factuality of the following sample > #Query#: <Here is the query for the LLM> #Response#: <Here is the response from the LLM according to the above query> #Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
Table 3: Instruction to generate evidence-based chains of facts for the “Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
@@
We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses. This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden factuality labels. Detailed prompts are in Table 7 through  10 of  §A.
@@ 4.4 Generate Chain of Evidence
Our benchmark evaluates the detectors’ ability to not only identify hallucinations but also to provide effective explanations. It necessitates ChatGPT’s generation of coherent golden evidence chains grounded in factual knowledge for substantiating judgments. Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts. Detailed prompts are presented in Tables 3 These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
@@ 4.5 Quality Filtering via Human Annotator
We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation. 21 educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting. To minimize subjectivity, we divided them into seven triads, introducing a voting mechanism for data evaluation. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively. We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations. We calculate $\kappa$ over a set of 100 annotated samples, resulting in a $\kappa$ value of 0.808, indicating substantial agreement.
@@
Zero-Shot
GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
@@
ICL (4-shot)
GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
@@
Det.(tune)
Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
@@
Knowledge
Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
@@ 5.2 Experimental Settings
We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT). Additionally, we explore the adoption of open-source models such as Llama2-chat et al (2023b), Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023), which are fine-tuned variants of the LLaMA Touvron et al. (2023).
@@
Using Azure’s OpenAI ChatGPT API [^3] https://openai.com/blog/chatgpt/, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation. For evaluations, we standardize the temperature at $0.2$ to minimize randomness. Detailed instructions are provided in Table 11 through Table 12 of §B.
@@ 5.3.2 In-context Learning Performance
The incorporation of few-shot information significantly improves fact-conflicting hallucination detection in the GPT-3.5-turbo, Alpaca-7B, Vicuna-7B, and Llama2-7B-chat models. Compared to models operating without this additional context, these enhancements result in an average increase of approximately $\uparrow$6in the FactCls and $\uparrow$18in ExpMatch scores. However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning. Furthermore, the improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of few-shot demonstrations.
@@ 5.3.4 Knowledge Enhancemant
Query: Who was the mother of Afonso II, the third king of Portugal?
Response:The mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile.
Correct Answer: Dulce Berenguer of Barcelona.
Returned by the tool: 1.—died March 25, 1223, Coimbra), the third king of Portugal (1211–23), under whom the reconquest of the south from the Muslims was continued. born: 1185?, Coimbra, Port. Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. 2…
Truth Guardian: FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct. This is supported by the evidence that Afonso II was the third king of Portugal and that Queen Urraca of Castile was his mother. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: NON-FACTUAL. The answer stating that the mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile is incorrect. According to the provided knowledge, the mother of Afonso II was Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, there is an incorrect conclusion in this question and answer.
@@
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict: NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
Query: Who is the Jeff Probst Show-nominated television producer’s first wife?
Response:His first wife was psychotherapist Shelley Wright.
Correct Answer: Shelley Wright.
Returned by the tool: 1. Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001. In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior. They broke up in early 2008. 2. …
Truth Guardian: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. According to the evidence provided, Jeff Probst was previously married to Shelley Wright, who is a *psychotherapist* in addition to being a television producer. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth Seeker: FACTUAL. According to the provided knowledge, Jeff Probst’s first wife was psychotherapist Shelley Wright. Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict:
@@
FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. Both AI models correctly identified that there are no fallacies, faulty reasoning, or incorrect conclusions present in the question and answer. The evidence provided confirms that Jeff Probst was married to Shelley Wright from 1996 to 2001, and that she is a psychotherapist in addition to being a television producer. Therefore, the answer is factual and accurate.

### v3

@@ Abstract
Despite their impressive generative capabilities, LLMs are hindered by fact-conflicting hallucinations in real-world applications. The accurate identification of hallucinations in texts generated by LLMs, especially in complex inferential scenarios, is a relatively unexplored area. To address this gap, we present FactCHD, a dedicated benchmark designed for the detection of fact-conflicting hallucinations from LLMs. FactCHD features a diverse dataset that spans various factuality patterns, including vanilla, multi-hop, comparison, and set operation. A distinctive element of FactCHD is its integration of fact-based evidence chains, significantly enhancing the depth of evaluating the detectors’ explanations. Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately. Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence.
@@ 1 Introduction
Large Language Models (LLMs) Zhao et al. (2023) are susceptible to generating text that, while seemingly credible, can be factually inaccurate or vague, leading to the spread of misinformation online Yin et al. (2023); Ji et al. (2023); Huang et al. (2023). This issue, referred to as fact-conflicting hallucination Zhang et al. (2023), arises from the incorporation of incorrect or obsolete knowledge into the models’ parameters and from the models’ inherent limitations in complex cognitive ability. These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information. Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
@@
However, traditional fact verification tasks Wadden et al. (2020a, b) are not suitable for LLM-based “Query-Response” data. Moreover, existing hallucination evaluation benchmarks Li et al. (2023); Muhlgay et al. (2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. To bridge this gap, we introduce an inherently rigorousWang et al. (2023b), yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
Datasets # Source Domains Factuality Pattern Evaluation Expandability
Fever Thorne et al. (2018) Text General Van. Consist.
Climate-Fever Diggelmann et al. (2020) Text Climate Change Van. Consist.
Health-Fever Sarrouti et al. (2021) Text Health Van. Consist.
Sci-Fact Wadden et al. (2020a) Text Scientific Van. Consist.
CoVERT Mohr et al. (2022) Text COVID-19 Van. Consist.
TabFact Chen et al. (2020) Table General Van. Consist.
HoVer Jiang et al. (2020) Text General Mul. Consist.
FEVEROUS Aly et al. (2021) Text+Table General Van. Consist.
HaluEval Li et al. (2023) Text General Van. Hallucination
FactCHD (ours) KGs ( [Uncaptioned image] ) & Text General&Vertical Van.&Mul.&Com.&Set. Fact.+Chain.
Table 1: Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the distinct factuality patterns.
@@
  • We present FactCHD [^1] Data is available at https://github.com/zjunlp/FactCHD., a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
@@
  • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses.
@@ 2 Related Work
Figure 3: Domain distribution of FactCHD, where “-p” and “-d” denote domains derived from Wikipedia and Wikidata, respectively.
Despite LLMs Zhao et al. (2023); Chen et al. (2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al. (2023); Wang et al. (2023a). These hallucinations, as categorized by Zhang et al. (2023), can be input-conflicting, context-conflicting, or fact-conflicting—with the latter being especially problematic due to the propagation of inaccurate factual information online. While previous studies have extensively examined hallucinations within natural language generation (NLG) for various NLP tasks Shuster et al. (2021); Creswell and Shanahan (2022); Das et al. (2023); Mallen et al. (2022), HaluEval Li et al. (2023) has emerged as a recent benchmark for assessing LLMs’ recognition of such errors. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate whether hallucinatory, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation.
@@
Our research is also related to prior works on fact verification in NLP tasks Thorne et al. (2018); Wadden et al. (2020b); Gupta et al. (2022); Dziri et al. (2022); Liang et al. (2023); Rashkin et al. (2021); Kryściński et al. (2020); Kang et al. (2024), expanding the understanding of factuality beyond binary judgments. The FRANK framework Pagnoni et al. (2021) offers a detailed typology of factual errors, while  Dhingra et al. (2019) explores lexical entailment in text generation. The FACTOR score by Muhlgay et al. (2023) evaluates LMs based on the likelihood of factual content. Existing benchmarks often miss the “Query-Response” context of LLMs, focusing solely on accuracy without providing explanatory rationales. Our contribution lies in: (1) presenting “Query-Response” formatted data for LLMs; (2) introducing metrics for interpretability in detecting fact-conflicting hallucinations; and (3) validating the effectiveness of Truth-Triangulator through cross-referential verification from multiple sources. The comparison with other datasets is detailed in Table 1.
@@ 4.1 Collect Realistic Data as Demonstration
Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM, Alpaca, and Vicuna) to generate authentic responses. By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence. Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
Figure 4: Overview of the construction process of FactCHD.
@@ 4.2 Generate “Query-Response” Contexts
We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses. This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden labels.
@@ 4.4 Generate Chain of Evidence
Our benchmark evaluates the detectors’ ability to not only identify hallucinations but also to provide effective explanations. It necessitates ChatGPT’s generation of coherent golden evidence chains grounded in factual knowledge for substantiating judgments. Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts. These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
@@ 4.5 Quality Filtering via Human Annotator
We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation. Several educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting. To minimize subjectivity, we organized them into groups of three, incorporating a voting mechanism for the evaluation of data. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively. We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations. We calculate $\kappa$ over the remaining annotated test set, resulting in a $\kappa$ value of 0.858, indicating substantial agreement.
@@
Zero-Shot GPT-3.5-turbo 55.12 22.79 59.54 29.84 16.66 18.89 55.46 28.23 52.82 24.03
@@
ICL (4-shot) GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
@@
Det.(tune) Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
@@
Knowledge Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
@@ 5.2 Experimental Settings
We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT). Additionally, we explore the adoption of open-source models such as Llama2-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat, AlpacaTaori et al. (2023) and Vicuna Chiang et al. (2023), which are fine-tuned variants of the LLaMA Touvron et al. (2023).
@@
Using Azure’s OpenAI ChatGPT API, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation. For evaluations, we standardize the temperature at $0.2$ to minimize randomness.
@@ 5.3.2 In-context Learning Performance
The incorporation of few-shot information significantly improves fact-conflicting hallucination detection in the GPT-3.5-turbo, Alpaca-7B, Vicuna-7B, and Llama2-7B-chat models. Compared to models operating without this additional context, these enhancements result in an average increase of approximately $\uparrow$6in the FactCls and $\uparrow$18in ExpMatch scores. However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning. The improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of demonstrations.
@@ 5.3.4 Knowledge Enhancemant
Query: Who was the mother of Afonso II, the third king of Portugal? Response:The mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile. Correct Answer: Dulce Berenguer of Barcelona. Returned by the tool: 1.—died March 25, 1223, Coimbra), the third king of Portugal (1211–23), under whom the reconquest of the south from the Muslims was continued. born: 1185?, Coimbra, Port. Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. 2… Truth Guardian: FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is correct. This is supported by the evidence that Afonso II was the third king of Portugal and that Queen Urraca of Castile was his mother. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer. Truth Seeker: NON-FACTUAL. The answer stating that the mother of Afonso II, the third king of Portugal, was Queen Urraca of Castile is incorrect. According to the provided knowledge, the mother of Afonso II was Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, there is an incorrect conclusion in this question and answer. Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict: NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
@@
Query: Who is the Jeff Probst Show-nominated television producer’s first wife? Response:His first wife was psychotherapist Shelley Wright. Correct Answer: Shelley Wright. Returned by the tool: 1. Probst was married to his first wife, psychotherapist Shelley Wright, from 1996 to 2001. In 2004, after filming on Survivor: Vanuatu wrapped, he began dating one of the contestants, a 23 year old Julie Berry - he was about 20 years her senior. They broke up in early 2008. 2. … Truth Guardian: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. According to the evidence provided, Jeff Probst was previously married to Shelley Wright, who is a *psychotherapist* in addition to being a television producer. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer. Truth Seeker: FACTUAL. According to the provided knowledge, Jeff Probst’s first wife was psychotherapist Shelley Wright. Therefore, the answer is correct and there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer. Truth-Triangulator: As the Fact Verdict Manager, after evaluating the responses from both AI models and examining the provided evidence, I have reached the following verdict: FACTUAL. The answer that Jeff Probst’s first wife was psychotherapist Shelley Wright is correct. Both AI models correctly identified that there are no fallacies, faulty reasoning, or incorrect conclusions present in the question and answer. The evidence provided confirms that Jeff Probst was married to Shelley Wright from 1996 to 2001, and that she is a psychotherapist in addition to being a television producer. Therefore, the answer is factual and accurate.